# Learning Optimal Stable Matches in Decentralized Markets with Unknown Preferences

### v1

@@ I Introduction
GS and similar algorithms are well-suited for situations where users report comprehensive preference orderings, but there are many applications where such requirements are prohibitive (Figure 1, bottom-left). Generating rankings may become difficult because of issues related to size, a lack of information, or both. For large practical markets like the National Resident Matching Program [6, 7, 11], several algorithms have been developed to handle the challenges associated with shortened or nonexistent preference orderings [12, 13].
@@
Our learning rule draws from the literature on learning in games, a well-established subfield of game theory concerned with the design of online learning algorithms [23, 24]. For several types of games, scholars have designed decentralized learning rules that guarantee convergence to desirable configurations such as correlated, Nash, and even welfare-maximizing equilibria [25, 26, 27, 28, 29]. Inspired by these results, we first formulate the decentralized two-sided matching market as a game in Section II. Then, in Section III, we assert the existence of a novel learning rule that guarantees convergence to the proposer-optimal stable match. We describe and simulate this rule in Sections IV and V, respectively, and prove its convergence properties in Section VII by considerably extending upon existing results from the literature on matching theory.
@@ II-A Market Model
The two-sided market consists of two groups: a set of proposers, denoted by $\mathbf{P}\triangleq\{P_{1},P_{2},\dots,P_{n}\}$, and a set of acceptors, denoted by $\mathbf{A}\triangleq\{A_{1},A_{2},\dots,A_{m}\}$. Each agent has a complete, transitive, and strict preference ordering over the agents in the other group. We represent the preference ordering of proposer $P_{i}$ as a function [^1] We choose this interval for simplicity, but any interval (open or closed) is acceptable $O_{P_{i}}:\mathbf{A}\cup\{\emptyset\}\to[0,1]$, where we say that $P_{i}$ prefers $A_{j_{1}}$ to $A_{j_{2}}$ if $O_{P_{i}}(A_{j_{1}})>O_{P_{i}}(A_{j_{2}})$, which we shorthand as $A_{j_{1}}\succ_{i}A_{j_{2}}$; for completeness, we assume $O_{P_{i}}(A_{j})>O_{P_{i}}(\emptyset)=0$ for all $j\in\{1,\dots m\}$. We define similarly the preference ordering $O_{A_{j}}$ of acceptor $A_{j}$, and as a slight abuse of notation, we use the same shorthand $\succ_{j}$ to describe the preferences of acceptor $A_{j}$. The collection of all preference orderings is given by $\mathbf{O}\triangleq\{O_{P_{1}},\dots,O_{P_{n}},O_{A_{1}},\dots,O_{A_{m}}\}$, so that a particular instance of the matching market game is fully parametrized by the tuple $\mathbf{M}=(\mathbf{P},\mathbf{A},\mathbf{O})$. A simple market is depicted in Figure 2 with $n=m=3$ proposers and acceptors.
@@ II-B Game Model
We denote the outcome of these acceptances and rejections as a match $\mu$. The proposers’ matches are described by $\{\mu_{P_{1}},\dots,\mu_{P_{n}}\}$, where $\mu_{P_{1}}(a)\in\mathbf{A}\cup\emptyset$ is the match (or lack thereof) of proposer $P_{i}$ given the action profile $a$; similarly, we write $\{\mu_{A_{1}},\dots,\mu_{A_{m}}\}$ for the acceptors where $\mu_{A_{j}}(a)\in\mathbf{P}\cup\emptyset$ denotes the outcome for $A_{j}$. If $\mu_{P_{i}}(a)=A_{j}$ then $\mu_{A_{j}}(a)=P_{i}$, and we say that the pair $(P_{i},A_{j})$ are matched. Otherwise, if $\mu_{P_{i}}(a)=\emptyset$ or $\mu_{A_{j}}(a)=\emptyset$, then we say that proposer $P_{i}$ or acceptor $A_{j}$ is unmatched.
@@
meaning that every proposer weakly prefers their partner in the POSM to their partner in any other stable match. For example, the proposer optimal stable match in the market shown in Figure 2 can be described by the list of proposer-acceptor pairs $\mu^{*}:(P_{1},A_{1}),(P_{2},A_{2}),(P_{3},A_{3})$.
From this, one can conclude that every pure strategy Nash equilibrium of the game corresponds to a stable matching. More specifically, every equilibrium action profile produces exactly one stable match; however, a stable match may be represented by multiple equilibrium action profiles if $n>m$, since the proposers who are unmatched in a stable match are indifferent between their actions. Furthermore, if one defines the welfare of an action profile as the sum of the players’ utilities corresponding to that action profile, then it is straightforward to show that any action profile corresponding to the proposer-optimal stable match is a welfare-maximizing equilibrium of the game.
@@ IV Description of Learning Rule
V Simulation
Figure 3: When using the learning rule, the empirical frequency of reaching any stable match (SM) and the POSM increases with the number of timesteps when every agent follows the learning rule with parameters $\epsilon=0.001$, $F(u)=-0.49\exp(-4u)$, and $G(\Delta u)=-0.49\exp(-4\Delta u)$.
Consider the market with three proposers and three acceptors shown in Figure 2. We simulate the performance of our learning rule on this market over 1,000 different initial matches for 100,000 timesteps. The empirical frequency of playing an action corresponding to a stable match or the POSM is shown for each timestep in the right portion of Figure 3. Observe that the empirical frequency with which the agents’ baseline action corresponds to the POSM increases with the number of periods, as does the frequency with which their action corresponds to any stable match.
We call attention to the fact that the empirical frequency of being at any stable match plateaus within the first 10,000 timesteps, while the frequency of being at the POSM increases. This is an important observation, as it shows that the dynamics successfully incentivize proposers to deviate from suboptimal stable matches. After escaping these configurations, they can then progress towards the POSM.
VI Conclusion
@@ VII Proof of Theorem
The proof of Theorem 1 relies on several existing results from the literature on learning in games and two-sided matching markets. First, in Sections VII-A and VII-B, we present and extend existing results from Markov process theory and matching theory, respectively. Then, we combine these results to complete our proof in Section VII-C.
VII-A Stochastic stability
Let $\mathbf{T}^{0}$ denote the probability transition matrix of a finite state Markov chain on a state space $\mathbf{X}$. Consider a ’perturbed’ version of the process, $\mathbf{T}^{\epsilon}$, where the ’size’ of the perturbation is described by a scalar $\epsilon>0$. We say that $\mathbf{T}^{\epsilon}$ is a regular perturbed Markov process if the following conditions hold:
  1. 1.
    $\mathbf{T}^{\epsilon}$ is ergodic
  2. 2.
    $\mathbf{T}^{\epsilon}$ approaches $\mathbf{T}^{0}$ at an exponentially smooth rate, meaning that if $\mathbf{T}^{\epsilon}_{xy}>0$ for some $x,y\in\mathbf{X}$ and for some $\epsilon>0$, then $0<\lim_{\epsilon\to 0^{+}}\frac{\mathbf{T}_{xy}^{\epsilon}}{\epsilon^{r(x,y)}}<\infty$.
We refer to the unique positive number $r(x,y)$ as the resistance of the transition from state $x$ to state $y$. For completeness, if $\mathbf{T}^{\epsilon}_{xy}=0$ for any $\epsilon>0$, then $r(x,y)=\infty$; similarly, if $\mathbf{T}^{0}_{xy}=0$, then $r(x,y)=0$.
Let $\mathbf{Z}\triangleq\{z_{0},z_{1},\dots,z_{v}\}$ denote the set of all recurrence classes of $\mathbf{T}^{0}$. For each pair of recurrence classes $z_{i},z_{j}$, an $ij$-path is a sequence of states $[x]=\{x_{1},x_{2},\dots,x_{k}\}$ such that $x_{1}\in z_{i}$ and $x_{k}\in z_{j}$. The resistance of the path $[x]$ is the sum of the resistances along each of its transitions, i.e., $r([x])=r(x_{1},x_{2})+\dots+r(x_{k-1},x_{k})$. We define the resistance from $z_{i}$ to $z_{j}$ as the minimum resistance over all $ij$-paths, which we denote as $\rho(z_{i},z_{j})$.
Now, consider a graph $G$ with $v$ vertices, each of which is a recurrence class of the unperturbed process $\mathbf{T}^{0}$, and let the weight of the edge from $z_{i}$ to $z_{j}$ be $\rho(z_{i},z_{j})$. We define a $z_{j}$-tree $T(z_{j})$ as a subgraph of $G$ such that $T(z_{j})$ has no cycles and for all $z_{i}\neq z_{j}$, there exists a unique path from $z_{i}$ to $z_{j}$ in $T(z_{j})$. The resistance of a $z_{j}$-tree is the sum of the resistances of the edges that compose it. The stochastic potential $\rho(z_{j})$ of a node $z_{j}$ is the minimum resistance over all $z_{j}$-trees. Lemma 1 provides a useful criterion for finding the stochastically stable states of a regular perturbed Markov process, which are those that are observed with positive probability in the limiting stationary distribution as $\epsilon\to 0$.
Lemma 1 (Young, 1993 [23]).
The stochastically stable states of a regular perturbed Markov process $\mathbf{T}^{\epsilon}$ are precisely the states contained in the recurrence classes of minimum stochastic potential.
VII-B Matching theory
Henceforth, we use the notation $\mu(x)$ to denote the partner of agent $x$ in the match $\mu$ for readability. In every two-sided market, the structure of the set of all stable matches can be described by a lattice [3]. Recall that a lattice $\mathbf{L}$ is a partially ordered set with the property that every pair of elements $x,y\in\mathbf{L}$ has a unique supremum and a unique infimum with respect to the set order $\leqslant$, both of which are contained in $\mathbf{L}$. For the lattice of stable matches, we define the order operation $\leqslant$ such that $\mu^{\prime}\leqslant\mu$ implies that every proposer weakly prefers their partner in $\mu$ to their partner in $\mu^{\prime}$.
Given a stable match $\mu\neq\mu^{*}$, we say that $\mu^{+}$ is a next-best stable match with respect to $\mu$ if $\mu^{+}$ lies directly above $\mu$ on the lattice of stable matches, i.e., there is no intermediate stable match $\mu^{\prime}$ such that $\mu\leqslant\mu^{\prime}\leqslant\mu^{+}$. Let $\{P_{i_{1}},\dots,P_{i_{r}}\}\subset\mathbf{P}$ denote the set of proposers who strictly prefer $\mu^{+}$ to $\mu$. Assuming that the indices $i_{1},\dots,i_{r}$ are ordered appropriately, we say that $\{P_{i_{1}},\dots,P_{i_{r}}\}$ form a rotation, meaning that
$\mu(P_{i_{k}})=\mu^{\prime}\left(P_{i_{k\ (\mathrm{mod}\ r)+1}}\right),\quad\forall k\in\{1,\dots,r\}.$
Hence, going between two adjacent stable matches simply requires cyclically shifting a subset of proposers. The concept of a rotation extends to arbitrary pairs of stable matches, but in this text, we exclusively use the term to describe a set of proposers who strictly improve from one stable match to another that lies directly above it on $\mathbf{L}$.
With these ideas in mind, we must design dynamics that can not only reach the lattice of stable matches, but also ’traverse’ it in an upwards manner. A useful tool for this task, termed best response dynamics, is introduced in [31]. Given an initial match $\mu^{0}$, the best response dynamics produce a sequence of matches in two phases:
  1. 1.
    During Phase 1, at each timestep, exactly one matched proposer updates their action to their best response, meaning that they propose to their most preferred acceptor who will accept them. This produces a finite sequence of matches $\mu^{0},\mu^{1},\dots,\mu^{k}$ with the property that in $\mu^{k}$, every matched proposer is playing a best response.
  2. 2.
    During Phase 2, at each timestep, exactly one unmatched proposer updates their action to their best response. This produces a finite sequence of matches $\mu^{k},\mu^{k+1},\dots,\mu^{l}$.
The best response dynamics satisfy the following property:
Lemma 2 (Ackermann et al., 2008, [31]).
For every matching market $\mathbf{M}$ with $n$ proposers and $m$ acceptors and every match $\mu^{0}$, there exists a sequence of at most $2nm$ best responses starting in $\mu^{0}$ that lead to a match $\mu^{l}$ that is stable.
Consider a stable match $\mu$, and define the match $\mu^{-i}$ as the near-stable match resulting from $\mu$ if proposer $P_{i}$ is made single, i.e.,
$\mu^{-i}(P_{j})=\begin{cases}\mu(P_{j})&j\neq i\\ P_{j}&j=i.\end{cases}$
The following result establishes a useful property of the best response dynamics initiated at near-stable matches.
Lemma 3.
Consider an arbitrary matching market $\mathbf{M}$ and an associated stable match $\mu\neq\mu^{*}$, and let $\mathbf{P}^{\prime}=\{P_{i_{1}},\dots,P_{i_{r}}\}\subseteq\mathbf{P}$ be a rotation for some next-best stable match with respect to $\mu$. For every $i\in\{i_{1},\dots i_{r}\}$, every best response sequence starting from the near-stable match $\mu^{-i}$ leads to a stable match $\mu^{\prime}>\mu$.
Before proving this result, we state a useful corollary:
Corollary 1.
Consider an arbitrary matching market $\mathbf{M}$ and an associated stable match $\mu$. For every $i\in\{1,\dots n\}$, every best response sequence starting from the near-stable match $\mu^{-i}$ leads to a stable match $\mu^{\prime}\geqslant\mu$.
We first prove Lemma 3, from which the proof of Corollary 1 follows.
Proof.
Let $\mu^{+}$ denote the next-best stable match from $\mu$ with respect to the rotation $\mathbf{P}^{\prime}$, let $\mu^{k}$ denote the (possibly unstable) match achieved at the end of Phase 1 in the best response dynamics, and let $\mu^{\prime}$ denote the stable match achieved at the end of Phase 2. The proof proceeds in two parts, each of which establishes a claim regarding Phase 1 and Phase 2, respectively.
First, we claim that every proposer who is matched at the end of Phase 1 weakly prefers their partner in $\mu^{k}$ to their partner in $\mu^{+}$. Suppose that proposer $P_{x}$ is matched in $\mu^{k}$, i.e., $\mu^{k}(P_{x})\neq P_{x}$. Recall that in Phase 1, only matched proposers play best responses; unmatched proposers do nothing. Thus, if $P_{x}$ is matched at the end of Phase 1, then at the least, they weakly prefer their partner in $\mu^{k}$ to their partner in $\mu^{-i}$. If $P_{x}\notin\mathbf{P}^{\prime}$, then the claim is certainly true, since in this case, $\mu^{+}(P_{x})=\mu^{-i}(P_{x})$. Suppose instead that $P_{x}\in\mathbf{P}^{\prime}$. We claim that if $\mu^{+}(P_{x})\succ_{j}\mu^{k}(P_{x})$, then at least one proposer is not playing a best response in $\mu^{k}$. Without loss of generality, let $P_{i_{1}}=P_{x}$ and let $A_{j_{1}}=\mu^{+}(P_{i_{1}})$ denote the partner of proposer $P_{i_{1}}$ in $\mu^{+}$. Consider the following three cases:
  • Case 1: $A_{j_{1}}$ is single or matched with someone they prefer less than $P_{i_{1}}$ in $\mu^{k}$. In this case, $P_{i_{1}}$ is not playing a best response, since they could propose to and match with $A_{i_{1}}$ instead, whom they prefer to their current partner.
  • Case 2: $A_{j_{1}}$ is matched with a proposer $P_{y}\notin\mathbf{P}^{\prime}$ whom they prefer to $P_{i_{1}}$. Since $P_{y}$ is not in the rotation, it must be the case that $P_{y}$ is not matched with $A_{j_{1}}$ in $\mu^{+}$. This implies that $P_{y}$ must strictly prefer $A_{j_{1}}$ to their partner in $\mu^{+}$, but this would imply that $\mu^{+}$ is unstable, which is a contradiction.
  • Case 3: $A_{j_{1}}$ is matched with a proposer $P_{i_{2}}\in\mathbf{P}^{\prime}$ whom they prefer to $P_{i_{1}}$. By the previous reasoning, $P_{i_{2}}$ must not prefer $A_{i_{1}}$ to their partner in $\mu^{+}$, whom we denote as $A_{j_{2}}$. The relationship between $P_{i_{2}}$ and $A_{j_{2}}$ must also fall in one of the three cases discussed presently. If it falls in Case 1, then $P_{i_{2}}$ is not playing a best response, and if it falls in Case 2, it would imply a contradiction. Suppose then that it falls into Case 3, so that $A_{j_{2}}$ prefers $P_{i_{3}}$, and generalize this idea so that the sequence of proposers $P_{i_{1}},P_{i_{2}},\dots,P_{i_{q}}$ fall into Case 3, where each $P_{i_{k}}$ seeks to be matched with $A_{i_{k}}$, their partner in $\mu^{+}$. Because the near-stable match $\mu^{-i}$ was obtained by making proposer $P_{i}\in\mathbf{P}^{\prime}$ single, we must have that $q$ is strictly less than $r$. However, by the pigeonhole principle, this would imply that two proposers are matched with the same acceptor in $\mu^{+}$, which is a contradiction.
These cases establish the claim that if $\mu^{+}(P_{i_{1}})\succ_{i_{1}}\mu^{k}(P_{i_{1}})$, then at least one proposer is not playing a best response in $\mu^{k}$. In turn, this establishes the first claim that every proposer who is matched at the end of Phase 1 weakly prefers their partner in $\mu^{k}$ to their partner in $\mu^{+}$.
Next, we discuss Phase 2. Consider a match $\mu^{k}$ in which every matched proposer weakly prefers their partner in $\mu^{k}$ to their partner in $\mu^{+}$. We claim that if an unmatched proposer $P_{x}$ plays a best response to $\mu^{k}$, generating the match $\mu^{k+1}$, then every matched proposer in $\mu^{k+1}$ will weakly prefer their partner in $\mu^{k+1}$ to their partner in $\mu^{+}$. This amounts to showing that $P_{x}$ can at least match with $A_{x}$, their partner in $\mu^{+}$. If $A_{x}$ is unmatched, then $P_{x}$ can certainly match with $A_{x}$; otherwise, suppose $A_{x}$ is matched with $P_{y}$. It must be the case that $A_{x}$ prefers $P_{x}$ to $P_{y}$; if this were not true, then $P_{y}$ and $A_{x}$ would both prefer one another to their partners in $\mu^{+}$, which would contradict the assumption that $\mu^{+}$ is stable. Thus, the second claim is established.
After having established these two claims, the proof of Lemma 3 is essentially complete. Phase 1 of the best response dynamics generates a match $\mu^{k}$ that every matched proposer weakly prefers to $\mu^{+}$, and every best response update performed by an unmatched proposer in Phase 2 preserves this property. The number of best response updates in each phase is finite, leading to a stable match which every proposer weakly prefers to $\mu^{+}$. Importantly, the proof is agnostic to the order in which the best response updates are executed within each Phase.
The proof of Corollary 1 follows from the fact that if $P_{i}\notin\mathbf{P}^{\prime}$, then Phase 1 will terminate at a match $\mu^{k}$ that every matched proposer weakly prefers to $\mu$, but not necessarily $\mu^{+}$. However, the previous arguments regarding Phase 2 are identical, so the best response dynamics will still terminate at a match that every proposer weakly prefers to $\mu$. ∎
VII-C Proof
@@
In this section, we combine the results from each of the previous sections to complete our proof of Theorem 1. Observe that the learning rule induces a regularly perturbed Markov process over the finite state space
@@
where $\text{ran}(f)$ denotes the range of the function $f$. We must show that the stochastically stable states of this process correspond to the POSM. We write the probability transition matrix over the state space $\mathbf{X}$ as $\mathbf{T}^{\epsilon}$ for each $\epsilon>0$. We must first identify the recurrence classes of the unperturbed process $\mathbf{T}^{0}$. Let $\mathbf{Z}\subseteq\mathbf{X}$ denote the subset of states in which every agents’ benchmark utilities are aligned with their benchmark actions, i.e., no agent is watchful. Let $\mathbf{C}$ denote the subset of states in which at least one matched (content) proposer is not playing a best response, and let $\mathbf{D}$ denote the subset of states in which every matched proposer is playing a best response, but there is at least one unmatched (discontent) proposer. Let $\mathbf{E}\subset\mathbf{Z}$ denote the subset of states whose benchmark action profiles correspond to stable matches, and let $e^{*}\in\mathbf{E}$ denote the state whose benchmark action profile correspond to the POSM. Observe that $\mathbf{C}$, $\mathbf{D}$, and $\mathbf{E}$ form a partition of $\mathbf{Z}$.
@@
Let $G$ be the graph constructed from $\mathbf{Z}$ as described in Section VII-A. We say that the edge from $z$ to $z^{\prime}$ is easy if $r(z,z^{\prime})=r^{*}(z)$. Similarly, we say that a path or tree is easy if each of its constituent edges are easy. The remainder of the proof involves showing that easy paths leads to the POSM, implying that the only easy tree is rooted at the POSM. We begin by identifying the nature of easy transitions for various states.
First, consider a state $z\in\mathbf{C}$, and suppose that the baseline action profile is such that at least one content proposer is not playing a best response. We claim that any easy edge from $z$ to some other state $z^{\prime}$ must involve a content proposer updating to a best response action. To see this, observe that the only easy transitions out of $z$ require experimentation. In particular, proposers can induce a transition out of $z$ by experimenting twice in succession without updating their baseline actions, causing at least one proposer to first become watchful and then discontent. Every other possible easy transition out of $z$ requires a proposer to experiment and update their baseline action: the resistance of a transition whereby a content proposer chooses to remain single is $2$, the resistance of a transition whereby a discontent proposer chooses to experiment and update their action is $1.5+F(\Delta u)$, and the resistance of a transition whereby a content proposer chooses to experiment and update their action is $1+G(\Delta u)$, where $\Delta u$ is the change in utility that the updating proposer experiences. Since $0.50\leq F(\cdot),G(\cdot)<$ 0.5, it is clear that a transition of minimum resistance must involve a content proposer experimenting and updating their action; moreover, since $G$ is strictly monotone decreasing, this update must be a best response. Given that no one experiments in the following timestep, which occurs with probability $O(1)$, the process transitions to a state $z^{\prime}\in\mathbf{Z}$ in two timesteps, and the total resistance along this path is $1+G(\Delta u)$.
@@
Having established the nature of these easy paths, we can now provide a simple construction of an easy tree rooted at $e^{*}$. Initialize $T(e^{*})$ as an empty set of edges, and add to it as follows. Consider a vertex $z_{1}\in\mathbf{Z}\setminus\mathbf{E}$. If $z_{1}$ is not part of an edge already in $T(e^{*})$, then add the easy edge from $z_{1}$ to $z_{2}$. Continue adding edges in this fashion, traversing nodes $z_{3},\dots,z_{k}$ until either $z_{k}$ already has an outgoing edge in $T(e^{*})$, or until reaching an equilibrium $z_{k}=e_{1}\in\mathbf{E}$. Repeat this process until every node in $\mathbf{Z}\setminus\mathbf{E}$ has exactly one outgoing edge. Then, for every vertex $e\in\mathbf{E}\setminus\{e^{*}\}$, add an easy edge from $e$ to a state $z\in\mathbf{Z}$ whose baseline action constitutes a near-stable match; in particular, choose a state such that the proposer made single belongs to a rotation with respect to the stable match represented by $e_{1}$. By Lemma 3, the easy path from $e$ will lead to an equilibrium that every proposer weakly prefers to a next-stable match with respect to the stable match given by $e$, because a proposer in a rotation is made single. It is straightforward to show that the properties of the best response dynamics ensure that this construction does not introduce any cycles. Thus, there exists a unique easy path from every vertex $z\in\mathbf{Z}\setminus\{e^{*}\}$ to $e^{*}$ in $T(e^{*})$, so $T(e^{*})$ is indeed an easy tree rooted at $e^{*}$.

### v2

@@ I Introduction
GS and similar algorithms are well-suited for situations where users report comprehensive preference orderings, but there are many applications where such requirements are prohibitive (Figure 1, bottom-left). Generating rankings may become difficult because of issues related to size, a lack of information, or both. To address these issues, large practical markets like the National Resident Matching Program [6, 7, 11] use specialized algorithms designed to handle shortened or nonexistent preference orderings [12, 13].
@@
Our learning rule draws from the literature on learning in games, a well-established subfield of game theory concerned with the design of online learning algorithms [23, 24]. For several types of games, scholars have designed decentralized learning rules that guarantee convergence to desirable configurations such as correlated, Nash, and even welfare-maximizing equilibria [25, 26, 27, 28, 29]. Inspired by these results, we first formulate the decentralized two-sided matching market as a game in Section II. Then, in Section III, we assert the existence of a novel learning rule that guarantees convergence to the proposer-optimal stable match. We describe this rule in Section IV and prove its convergence properties in Section VI. Simulations and additional proof details are available in the full version of the text [30].
@@ II-A Market Model
The two-sided market consists of two groups: a set of proposers, denoted by $\mathbf{P}\triangleq\{P_{1},P_{2},\dots,P_{n}\}$, and a disjoint iset of acceptors, denoted by $\mathbf{A}\triangleq\{A_{1},A_{2},\dots,A_{m}\}$. Each agent has a complete, transitive, and strict preference ordering, also referred to as preferences, over the agents in the other group. We represent the preference ordering of proposer $P_{i}$ as a function [^1] We choose this interval for simplicity, but any bounded interval (open or closed) is acceptable. $O_{P_{i}}:\mathbf{A}\cup\{\emptyset\}\to[0,1]$, where we say that $P_{i}$ prefers $A_{j_{1}}$ to $A_{j_{2}}$ if $O_{P_{i}}(A_{j_{1}})>O_{P_{i}}(A_{j_{2}})$, which we shorthand as $A_{j_{1}}\succ_{i}A_{j_{2}}$; for completeness, we assume $O_{P_{i}}(A_{j})>O_{P_{i}}(\emptyset)=0$ for all $j\in\{1,\dots m\}$. We define similarly the preference ordering $O_{A_{j}}$ of acceptor $A_{j}$, and as a slight abuse of notation, we use the same shorthand $\succ_{j}$ to describe the preferences of acceptor $A_{j}$. The collection of all preference orderings is given by $\mathbf{O}\triangleq\{O_{P_{1}},\dots,O_{P_{n}},O_{A_{1}},\dots,O_{A_{m}}\}$, so that a particular instance of the matching market game is fully parametrized by the tuple $\mathbf{M}=(\mathbf{P},\mathbf{A},\mathbf{O})$. A simple market is depicted in Figure 2 with $n=m=3$ proposers and acceptors.
@@ II-B Game Model
We denote the outcome of these acceptances and rejections as a match $\mu(a)\triangleq\{\mu_{P_{1}}(a),\dots,\mu_{P_{n}}(a),\mu_{A_{1}}(a),\dots,\mu_{A_{m}}(a)\}$, where $\mu_{P_{i}}(a)\in\mathbf{A}\cup\{\emptyset\}$ is the match (or lack thereof) of proposer $P_{i}$ given the action profile $a$, and $\mu_{A_{j}}(a)\in\mathbf{P}\cup\{\emptyset\}$ denotes the outcome for $A_{j}$. If $\mu_{P_{i}}(a)=A_{j}$ then $\mu_{A_{j}}(a)=P_{i}$, and we say that the pair $(P_{i},A_{j})$ are matched. Otherwise, if $\mu_{P_{i}}(a)=\emptyset$ or $\mu_{A_{j}}(a)=\emptyset$, then we say that proposer $P_{i}$ or acceptor $A_{j}$ is unmatched. For ease of notation, we simply write $\mu$, $\mu_{P_{i}}$ or $\mu_{A_{j}}$ when the dependence on $a$ is clear.
@@
meaning that every proposer weakly prefers their partner in the POSM to their partner in any other stable match. For example, the proposer optimal stable match in the market shown in Figure 2 can be described by the list of proposer-acceptor pairs $(P_{1},A_{1}),(P_{2},A_{2}),(P_{3},A_{3})$.
From this, one can conclude that every pure strategy Nash equilibrium of the game corresponds to a stable matching. More specifically, every Nash equilibrium action profile produces exactly one stable match; however, a stable match may be represented by multiple Nash equilibrium action profiles if $n>m$, since the proposers who are unmatched in a stable match are indifferent between their actions. Furthermore, if one defines the welfare of an action profile as the sum of the players’ utilities corresponding to that action profile, then it is straightforward to show that any action profile corresponding to the proposer-optimal stable match is a welfare-maximizing Nash equilibrium of the game.
@@ IV Description of Learning Rule
V Conclusion
@@ VI Proof of Theorem
The proof of Theorem 1 relies on several existing results from the literature on learning in games and two-sided matching markets, as well as some novel results on best-response dynamics in matching games available in the full version of the text [30]. We combine these results to complete our proof below.
@@
Observe that the learning rule induces a regularly perturbed Markov process over the finite state space
@@
where $\text{ran}(f)$ denotes the range of the function $f$. We must show that the stochastically stable states of this process correspond to the POSM [23, Theorem 4] We write the probability transition matrix over the state space $\mathbf{X}$ as $\mathbf{T}^{\epsilon}$ for each $\epsilon>0$. We must first identify the recurrence classes of the unperturbed process $\mathbf{T}^{0}$. Let $\mathbf{Z}\subseteq\mathbf{X}$ denote the subset of states in which every agents’ benchmark utilities are aligned with their benchmark actions, i.e., no agent is watchful. Let $\mathbf{C}$ denote the subset of states in which at least one matched (content) proposer is not playing a best response, and let $\mathbf{D}$ denote the subset of states in which every matched proposer is playing a best response, but there is at least one unmatched (discontent) proposer. Let $\mathbf{E}\subset\mathbf{Z}$ denote the subset of states whose benchmark action profiles correspond to stable matches, and let $e^{*}\in\mathbf{E}$ denote the state whose benchmark action profile correspond to the POSM. Observe that $\mathbf{C}$, $\mathbf{D}$, and $\mathbf{E}$ form a partition of $\mathbf{Z}$.
@@
Let $G$ be the graph constructed from $\mathbf{Z}$ as described in Section VII-A in [30]. We say that the edge from $z$ to $z^{\prime}$ is easy if $r(z,z^{\prime})=r^{*}(z)$. Similarly, we say that a path or tree is easy if each of its constituent edges are easy. The remainder of the proof involves showing that easy paths leads to the POSM, implying that the only easy tree is rooted at the POSM. We begin by identifying the nature of easy transitions for various states.
First, consider a state $z\in\mathbf{C}$, and suppose that the baseline action profile is such that at least one content proposer is not playing a best response. We claim that any easy edge from $z$ to some other state $z^{\prime}$ must involve a content proposer updating to a best response action. To see this, observe that the only possible easy transitions out of $z$ require experimentation. In particular, proposers can induce a transition out of $z$ by experimenting twice in succession without updating their baseline actions, causing at least one proposer to first become watchful and then discontent. Every other possible easy transition out of $z$ requires a proposer to experiment and update their baseline action: the resistance of a transition whereby a content proposer chooses to remain single is $2$, the resistance of a transition whereby a discontent proposer chooses to experiment and update their action is $1.5+F(\Delta u)$, and the resistance of a transition whereby a content proposer chooses to experiment and update their action is $1+G(\Delta u)$, where $\Delta u$ is the change in utility that the updating proposer experiences. Since $0.50\leq F(\cdot),G(\cdot)<$ 0.5, it is clear that a transition of minimum resistance must involve a content proposer experimenting and updating their action; moreover, since $G$ is strictly monotone decreasing, this update must be a best response. Given that no one experiments in the following timestep, which occurs with probability $O(1)$, the process transitions to a state $z^{\prime}\in\mathbf{Z}$ in two timesteps, and the total resistance along this path is $1+G(\Delta u)$.
@@
Having established the nature of these easy paths, we can now provide a simple construction of an easy tree rooted at $e^{*}$. Initialize $T(e^{*})$ as an empty set of edges, and add to it as follows. Consider a vertex $z_{1}\in\mathbf{Z}\setminus\mathbf{E}$. If $z_{1}$ is not part of an edge already in $T(e^{*})$, then add the easy edge from $z_{1}$ to $z_{2}$. Continue adding edges in this fashion, traversing nodes $z_{3},\dots,z_{k}$ until either $z_{k}$ already has an outgoing edge in $T(e^{*})$, or until reaching an equilibrium $z_{k}=e_{1}\in\mathbf{E}$. Repeat this process until every node in $\mathbf{Z}\setminus\mathbf{E}$ has exactly one outgoing edge. Then, for every vertex $e\in\mathbf{E}\setminus\{e^{*}\}$, add an easy edge from $e$ to a state $z\in\mathbf{Z}$ whose baseline action constitutes a near-stable match; in particular, choose a state such that the proposer made single belongs to a rotation with respect to the stable match represented by $e_{1}$. However, from [30, Lemma 3], the easy path from $e$ will lead to an equilibrium that every proposer weakly prefers to a next-stable match with respect to the stable match given by $e$, because a proposer in a rotation is made single. It is straightforward to show that the properties of the best response dynamics ensure that this construction does not introduce any cycles. Thus, there exists a unique easy path from every vertex $z\in\mathbf{Z}\setminus\{e^{*}\}$ to $e^{*}$ in $T(e^{*})$, so $T(e^{*})$ is indeed an easy tree rooted at $e^{*}$.